OpenStudy (anonymous):

show work in substitution/elimination thank (1) x=y-2 4x+y=2s (2) y=3x-2 y=2x-5 (3) -4x+5y=17 4x=6y=-6 (4) 8x+3y=4 -7x+5y=-34 (5) 6x+y=-39 3x+2y=-15

OpenStudy (anonymous):

I will do number 2. I believe you can just add the equations, to give you y=5x+7. but I am not sure.

OpenStudy (anonymous):

I messaged you about number one, here are instructions for two and three. Try four and five on your own, but if you have trouble I will help. 2.) In this equation, y is already isolated, so we just plug in the isolated equation to the other equation. So y=3x-2 if and y=2x-3. than 3x-2=2x-3. now just solve for x x=-1 Now that we know x, we can plug it into either equation and solve for y y=3(-1)-2 y=-5 3.) -4x+5y=17 4x+6y=-6 We can isolate x in the second equation by subtracting x and dividing by 4 x=-1.5y-1.5 We can plug in the other equation -4(-1.5y-1.5)+5y=17 Then solve for y by distribution 6y+6+5y=17 combine like terms 11y+6=17subtract six 11y=11divide by 11 y=1 Now we plug y into one of the equations and solve for x -4x+5(1)=17 -4x=12 x=-3